# TurboMap: GPU-Accelerated Local Mapping for Visual SLAM

### v1

@@ IV-F Local Keyframe Culling
Local Keyframe Culling is the final major component of local mapping. TurboMap does not offload this stage to the GPU, mainly due to the significant data transfer overhead it would have. Instead, we apply a CPU-side optimization to improve its performance while avoiding the cost of GPU data transfers.
As described in Section III, the Keyframe Culling stage removes redundant keyframes by checking how many of their associated map points are also observed by other keyframes in the map at the same or finer scale, and discarding those that exceed a predefined threshold. To perform this check, the system maintains an observation list for each map point, which records the set of keyframes that observe that point. This data structure is large and frequently updated during each iteration of local mapping. Unlike the relatively static keyframe data which can be stored on the GPU without frequent synchronization, the dynamic nature of the observation data makes it unsuitable for the same treatment. We experimented with a GPU-side map point storage system that continuously updated this data, but found that maintaining consistency introduced significant overhead. In fact, the cost of keeping this storage synchronized was nearly equivalent to the total cost of the entire local mapping stage, making the approach impractical. As a result, we chose not to offload this component to the GPU.
Instead of offloading this logic to the GPU, we optimize it on the CPU by introducing a lightweight array that tracks how many times each map point is observed at each scale. Each time a map point is observed, the corresponding counter in the array is incremented. This approach eliminates the need to iterate over every keyframe’s observations to count the number of valid observations per map point. As a result, it significantly reduces the computational overhead of the keyframe culling process. The additional memory overhead of this optimization is minimal, as we store only a small array per map point, which amounts to just a few bytes per map point.
Sequence Total Local Mapping Speed Triangulation Speed Fusion Speed LBA Speed KF Culling Speed ATE
@@
Sequence Total Local Mapping Speed Triangulation Speed Fusion Speed LBA Speed KF Culling Speed ATE
@@ V-A Experimental Setup
We evaluate the performance of TurboMap by running a set of sequences on two different hardware platforms and comparing the results with those of the original ORB-SLAM3. The experiments are conducted in two setups: a desktop machine and a mobile embedded system. The desktop system is equipped with a 20-core Intel Core i7-12700K CPU running at 5.0 GHz, an NVIDIA RTX 3090 GPU with 10,496 CUDA cores, and 64 GB of RAM. The embedded platform is a Jetson Xavier NX, featuring a 6-core ARM Carmel CPU running at 1.4 GHz (20W configuration), a 384-core NVIDIA Volta GPU, and 8 GB of RAM. To ensure consistent and reproducible results, we lock the CPU and GPU frequencies to their maximum values during all experiments.
For evaluation, we use a combination of sequences from the TUM-VI and EuRoC datasets, both in the stereo-inertial configuration. We run each sequence five times and report the average results across these runs to account for variability and ensure fair comparison.
@@ V-B Overall Performance
Table I presents the performance comparison between TurboMap and ORB-SLAM3 in different components of local mapping. As shown, TurboMap gets an average speedup of $1.3\times$ in EuRoC and $1.6\times$ in TUM-VI, in both the desktop and embedded settings. Importantly, TurboMap achieves these performance gains while maintaining a similar Absolute Trajectory Error (ATE) RMSE as ORB-SLAM3, demonstrating that our optimizations do not compromise the accuracy of the system.
@@
To illustrate the impact of reduced variance, Figure 4 shows the keyframe processing time for both systems in the corridor1 sequence. As the number of keyframes in the map increases, ORB-SLAM3 exhibits a noticeable growth in processing time per keyframe. This is due to the increased workload across components such as triangulation, fusion, and local bundle adjustment, all of which scale with map size. In contrast, TurboMap maintains stable processing times even as the map grows, thanks to its GPU-accelerated components and CPU-side optimizations that effectively distribute and manage the computational load. This effect is clearly visible when comparing the maximum keyframe processing times in both systems: approximately 220 ms for ORB-SLAM3 versus only 115 ms for TurboMap.
@@ V-C Individual Component Performances
We analyze the performance gains from each optimization introduced in TurboMap in the following subsections. It is important to note that the observed speedups differ between the EuRoC and TUM-VI datasets. This discrepancy arises from the different camera models and the way ORB-SLAM3 processes them. Specifically, in TUM-VI, which employs fisheye cameras, the system does not rectify the images. Instead, it treats the stereo pair as two independent monocular cameras in order to leverage the wider field of view and retain more detected features. As a result, TUM-VI sequences contain approximately twice as many map points as those in EuRoC. Consequently, local mapping components, particularly those that scale with the number of map points, require significantly more processing time in TUM-VI.
V-C1 Search for Triangulation
The GPU optimization of the Search for Triangulation component results in an average speedup of $2.2\times$ on the EuRoC dataset and $3.3\times$ on the TUM-VI dataset in the desktop setting. In the Jetson setting, it achieves an average speedup of $3.5\times$ on EuRoC and $5.2\times$ on TUM-VI. The higher speedup achieved in TUM-VI is because the system is searching for more map point matches to triangulate in this dataset as we explained earlier, and TurboMap uses more GPU threads in this optimization compared to EuRoC, resulting in greater performance gains.
@@ V-C2 Map Point Fusion
Offloading the Map Point Fusion component to the GPU yields an average speedup of $1.5\times$ on the EuRoC dataset and $2.3\times$ on the TUM-VI dataset in the desktop setting. On the Jetson platform, it achieves average speedups of $1.4\times$ and $2.2\times$ on EuRoC and TUM-VI, respectively. The improved performance on TUM-VI is attributed to the higher number of map points being fused, allowing TurboMap to exploit parallelism more effectively by utilizing a greater number of GPU threads.
@@ V-C3 Keyframe Culling
TurboMap optimizes the Keyframe Culling process on the CPU, achieving an average speedup of $1.8\times$ on the EuRoC dataset in both desktop and Jetson settings. In the TUM-VI dataset, it achieves a speedup of $3.3\times$ in the desktop setting and $3.2\times$ in Jetson.
Figure 5: Trajectory comparison between TurboMap and ORB-SLAM3 in corridor1.
V-D Trajectory
@@
ATE
#Skipped LBA
& KF Culling
ATE
#Skipped LBA
& KF Culling
@@
TABLE II: Comparison of TurboMap and ORB-SLAM3 under high keyframe insertion, showing ATE RMSE (m) and the number of skipped Local Bundle Adjustments/Keyframe Cullings across sequences in the desktop setting, averaged over 5 runs.
@@
V-E Stress Testing Local Mapping
@@
As shown in Table II, TurboMap significantly outperforms ORB-SLAM3 under heavy load. While ORB-SLAM3 produces very high absolute trajectory errors (ATE), TurboMap maintains an accuracy comparable to normal load conditions, which is substantially lower than the error in ORB-SLAM3. This difference is also evident in Figure 6, where the trajectory produced by ORB-SLAM3 in the MH01 sequence diverges significantly from the ground truth, while TurboMap produces a trajectory that closely aligns with it. This discrepancy can be explained by the way ORB-SLAM3 manages new keyframes. The local mapping module maintains a fixed-size keyframe queue and processes keyframes sequentially. When the queue is not empty, ORB-SLAM3 skips Local Bundle Adjustment and Keyframe Culling stages for the current keyframe and starts processing the next keyframe in the queue in order to save time and avoid keyframe drops. However, skipping these stages directly impacts trajectory accuracy. Table II further illustrates this point by reporting the number of times LBA and Keyframe Culling are skipped under heavy load in both systems. The number of skips is substantially higher in ORB-SLAM3, which is the primary factor contributing to its decreased accuracy.
@@ VI Conclusion
In this work, we present TurboMap, a GPU-accelerated and CPU-optimized backend for ORB-SLAM3 designed to address key performance bottlenecks in the local mapping process. TurboMap leverages GPU parallelism to accelerate the Search for Triangulation and Map Point Fusion stages, introduces an efficient CPU-side optimization for Keyframe Culling, incorporates an existing GPU-accelerated implementation of Local Bundle Adjustment, and minimizes data transfer overhead by employing GPU-resident keyframe storage. Collectively, these enhancements yield an average speedup of $1.3\times$ in the EuRoC dataset and $1.6\times$ in the TUM-VI dataset in local mapping. One limitation of the proposed approach is the overhead associated with transferring data to GPU memory, which can introduce latency. However, on platforms such as NVIDIA Jetson, where the CPU and GPU share a unified memory architecture, this overhead is largely eliminated. This highlights the potential for even greater performance gains on embedded platforms where memory sharing allows for more efficient parallel computation without the cost of data migration.

### v2

@@ IV-F Local Keyframe Culling
Local Keyframe Culling is the final major component of local mapping. We do not parallelize this stage, mainly due to the significant data transfer overhead it would have. Instead, we apply a CPU-side optimization to improve its performance while avoiding the cost of GPU data transfers.
As described in Section III, the Keyframe Culling stage removes redundant keyframes by checking how many of their associated map points are also observed by other keyframes in the map at the same or finer scale, and discarding those that exceed a predefined threshold. To perform this check, the system maintains an observation list for each map point, which records the set of keyframes that observe that point. This data structure is large and frequently updated during each iteration of local mapping. Unlike the relatively static keyframe data which can be stored on the GPU without frequent synchronization, the dynamic nature of the observation data makes it unsuitable for the same treatment. We experimented with a GPU-side map point storage system that continuously updated these data, but found that maintaining consistency introduced significant overhead. In fact, the cost of keeping this storage synchronized was nearly equivalent to the total cost of the entire local mapping stage, making the approach impractical. As a result, we chose not to parallelize this component on the GPU.
Instead of parallelizing this logic on the GPU, we optimize it on the CPU by introducing a lightweight array that tracks how many times each map point is observed at each scale. Each time a map point is observed, the corresponding counter in the array is incremented. This approach eliminates the need to iterate over every keyframe’s observations to count the number of valid observations per map point. As a result, it significantly reduces the computational overhead of the keyframe culling process. The additional memory overhead of this optimization is minimal, as we store only a small array per map point, which amounts to just a few bytes.
Sequence Total Local Mapping Speed KP Search Speed MP Fusion Speed LBA Speed KF Culling Speed ATE (m)
@@
Sequence Total Local Mapping Speed KP Search Speed MP Fusion Speed LBA Speed KF Culling Speed ATE (m)
@@ V-A Experimental Setup
We evaluate the performance of TurboMap by running a set of sequences on two different hardware platforms and comparing the results with those of the original ORB-SLAM3. The experiments are conducted in two setups: a desktop machine and a mobile embedded system. The desktop system is equipped with a 20-core Intel Core i7-12700K CPU running at 5.0 GHz, an NVIDIA RTX 3090 GPU with 10,496 CUDA cores, and 64 GB of RAM. The embedded platform is a Jetson Xavier NX, featuring a 6-core ARM Carmel CPU running at 1.4 GHz, a 384-core NVIDIA Volta GPU, and 8 GB of RAM. To ensure consistent and reproducible results, we lock the CPU and GPU frequencies to their maximum values during all experiments.
We use a combination of sequences from the TUM-VI and EuRoC datasets, both in the stereo-inertial configuration. We run each sequence five times and report the average results across these runs.
@@ V-B Overall Performance
Table I presents the performance comparison between TurboMap and ORB-SLAM3 for local mapping. The reported timings for each component include CPU–GPU data transfer overhead. We also conduct an ablation study where each accelerated component is enabled independently. The results show that each component operates modularly, and activating any individual optimization yields consistent improvements. As shown, TurboMap gets an average speedup of $1.3\times$ in EuRoC and $1.6\times$ in TUM-VI, in both the desktop and embedded settings, when all optimizations are activated. Importantly, TurboMap achieves these performance gains while maintaining a similar Absolute Trajectory Error (ATE) RMSE as ORB-SLAM3, demonstrating that our optimizations do not compromise the accuracy of the system.
@@
To illustrate the impact of reduced variance, Figure 4 shows the keyframe processing time for both systems in the Corridor1 sequence. As the number of keyframes in the map increases, ORB-SLAM3 exhibits a noticeable growth in processing time per keyframe. This is due to the increased workload across components such as triangulation, fusion, and local bundle adjustment, all of which scale with map size. In contrast, TurboMap maintains more stable processing times even as the map grows, thanks to its GPU-accelerated components and CPU-side optimizations that effectively distribute and manage the computational load. This effect is clearly visible when comparing the maximum keyframe processing times in both systems: approximately 220 ms for ORB-SLAM3 versus only 115 ms for TurboMap.
@@ V-C Individual Component Performances
We analyze the performance gains from each optimization introduced in TurboMap in the following subsections. It is important to note that the observed speedups differ between the EuRoC and TUM-VI datasets. This discrepancy arises from the different camera models and the way they are processed. Specifically, in TUM-VI, which employs fisheye cameras, the system does not rectify the images. Instead, it treats the stereo pair as two independent monocular cameras in order to leverage the wider field of view and retain more detected features. As a result, TUM-VI sequences contain approximately twice as many map points as those in EuRoC. Consequently, local mapping components, particularly those that scale with the number of map points, require significantly more processing time in TUM-VI.
V-C1 Keypoint Correspondence Search
The GPU parallelization of the Keypoint Correspondence Search component results in an average speedup of $2.2\times$ on EuRoC and $3.3\times$ on TUM-VI in the desktop setting. In the Jetson setting, it achieves an average speedup of $3.5\times$ on EuRoC and $5.2\times$ on TUM-VI. The higher speedup achieved in TUM-VI is because the system is searching for more keypoint matches to triangulate in this dataset as we explained earlier, and TurboMap uses more GPU threads in this optimization compared to EuRoC, resulting in greater performance gains.
@@ V-C2 Map Point Fusion
Parallelizing Map Point Fusion yields an average speedup of $1.5\times$ on EuRoC and $2.3\times$ on TUM-VI in the desktop setting. On the Jetson platform, it achieves average speedups of $1.4\times$ and $2.2\times$ on EuRoC and TUM-VI, respectively. The improved performance on TUM-VI is attributed to the higher number of map points being fused, allowing TurboMap to exploit parallelism more effectively with a greater number of GPU threads.
@@ V-C3 Keyframe Culling
TurboMap optimizes the Keyframe Culling process on the CPU, achieving a speedup of $1.8\times$ on EuRoC on both desktop and Jetson. In TUM-VI, it achieves a speedup of $3.3\times$ on desktop and $3.2\times$ on Jetson.
Figure 5: Timing breakdown of Map Point Creation and Fusion stages, with keyframe storage, in TurboMap on Corridor1.
V-D Map Point Creation and Fusion Timing Analysis
Figure 5 shows the average execution time of different stages of Map Point Creation and Fusion in TurboMap on Corridor1, evaluated on Desktop and Jetson platforms, along with Keyframe Storage timings. CPU execution time corresponds to the portions remaining on the CPU due to negligible latency or unavoidable synchronization. Data preparation time reflects the cost of converting CPU-resident data into GPU-compatible formats and vice versa. This overhead is higher in Fusion, since selected map points must be restructured before transfer, whereas Map Point Creation primarily uses keyframe data already stored on the GPU.
In Map Point Creation, kernel execution dominates, indicating a computation-bound workload. In Fusion, data preparation dominates, reflecting the cost of restructuring map point data. Although transferring keyframes to the GPU is the most expensive data movement step, it is performed only once per local mapping iteration, preventing repeated transfer overhead. Importantly, the size of CPU–GPU data transfers does not grow with the size of the map. The only data structure that grows with the map is the set of keyframes. In our design, keyframes are transferred to the GPU one at a time as they are created and then kept resident, avoiding repeated bulk transfers. As a result, the per-iteration data transfer cost remains stable as the map expands.
Figure 6: Trajectories: TurboMap vs. ORB-SLAM3 (Corridor1)
V-E Trajectory
@@
ATE #Skipped LBA & KF Culling ATE #Skipped LBA & KF Culling
@@
TABLE II: Comparison of TurboMap and ORB-SLAM3 with high keyframe insertion: ATE RMSE (m), skipped LBAs, and keyframe culling (avg. 5 runs, desktop setting).
@@
V-F Stress Testing Local Mapping
@@
As shown in Table II, TurboMap significantly outperforms ORB-SLAM3 under heavy load. While ORB-SLAM3 produces very high trajectory errors, TurboMap maintains an accuracy comparable to normal load conditions, which is substantially lower than the error in ORB-SLAM3. This difference is also evident in Figure 7, where the trajectory produced by ORB-SLAM3 in the MH01 sequence diverges significantly from the ground truth, while TurboMap produces a trajectory that closely aligns with it. This discrepancy can be explained by the way ORB-SLAM3 manages new keyframes. Local mapping maintains a fixed-size keyframe queue and processes keyframes sequentially. When the queue is not empty, ORB-SLAM3 skips LBA and Keyframe Culling stages for the current keyframe and starts processing the next keyframe in the queue in order to save time and avoid keyframe drops. However, skipping these stages impacts trajectory accuracy. Table II further illustrates this point by reporting the number of times LBA and Keyframe Culling are skipped under heavy load in both systems. The number of skips is substantially higher in ORB-SLAM3, which is the primary factor contributing to its decreased accuracy.
@@ VI Conclusion
In this work, we presented TurboMap, a GPU-accelerated and CPU-optimized backend for ORB-SLAM3 that addresses the primary performance bottlenecks in the local mapping pipeline. Unlike prior work that mainly accelerates bundle adjustment, TurboMap targets the remaining dominant components that are constrained by shared-state updates and synchronization. We restructure Map Point Creation to enable parallel Keypoint Correspondence Search on the GPU, redesign Map Point Fusion for efficient parallel execution, and optimize Keyframe Culling on the CPU. To support these GPU-parallelized stages efficiently, we introduce a persistent GPU-resident keyframe storage that minimizes data transfer and synchronization overhead. Together with the integration of an existing GPU-accelerated LBA solver, these changes enable efficient parallel execution while preserving correctness and system behavior. Experimental results show that TurboMap achieves an average local mapping speedup of $1.3\times$ on the EuRoC dataset and $1.6\times$ on the TUM-VI dataset without degrading accuracy. These results demonstrate that holistically restructuring local mapping and carefully managing data movement can substantially improve the performance of modern visual SLAM systems.